OpenStudy (anonymous):

Describe the meaning of the rate. 6 days each month A. Something happens 6 days per month. B. Each month is 6 days long. C. Each day is one month long. D. A day happens six times each month.

OpenStudy (anonymous):

@Love_Ranaa @King.Void. I think its A o.o

OpenStudy (solomonzelman):

I will give you an example, do you mind?

OpenStudy (anonymous):

No :] @SolomonZelman

OpenStudy (solomonzelman):

For example: `Sam goes to "Pita-King" (some pizza place) and spend $10 there, (once) every week.` That means the "rate is" `$10 per week` is the rate. (to be more precise "$10 spend on pizza every week" )

OpenStudy (solomonzelman):

lets go a little different. a person buys 10 pizzas every month.

OpenStudy (solomonzelman):

that means a rate is \(\large\color{black}{ \displaystyle 10 }\) pizzas \(\large\color{black}{ \displaystyle ~{\rm per~~month } }\).

OpenStudy (solomonzelman):

What if we didn't know what he is doing 10 times per month ? (if we didn't know that he is buying pizza 10 times per month, but instead does something else, which is not known to us)

OpenStudy (solomonzelman):

Then we would say the rate [just] is: \(\large\color{black}{ \displaystyle 10~ ~per~{\rm~month } }\)

OpenStudy (solomonzelman):

which would mean that "something is happening 10 times each month"

OpenStudy (solomonzelman):

(again, we don't know what this "something" is)

OpenStudy (anonymous):

So it would be D?.0.

OpenStudy (solomonzelman):

not exactly. 1 day happens 6 times a month? but we said that \(\normalsize\color{black}{ \displaystyle\underline{{\rm~"something"~~happens }} }\) 6 times/days per month.

OpenStudy (solomonzelman):

the "day" doesn't happen. What happens is the [unknown] "something".

OpenStudy (anonymous):

Something is only in A. Like I said .0. at the top

OpenStudy (anonymous):

@SolomonZelman Thanks for explaining the answer that I said it was :) so am I correct?

OpenStudy (solomonzelman):

yes, A is right.

OpenStudy (solomonzelman):

Also, you can exclude B and C just by reading what they say.

OpenStudy (solomonzelman):

month is never 6 days long, and certainly 1 day is not a month long.

OpenStudy (anonymous):

That was I was thinking!! @SolomonZelman

OpenStudy (solomonzelman):

yes:)